# Sorting inversion sequences

### v5

@@ 5.   Open Problems
And the cases restricting pop stacks to particular depths in an analogous way to  [14, 15, 16, 17] and/or the expansion in allowable pushes analogous to  [2] can be interesting for words and inversion sequences.
@@
Another avenue of exploration is to consider other sorting machines or algorithms. For instance, the authors considered a sorting algorithm on a single stack prioritizing outputting the next correct entry of the sequence until no other moves were available and then either returning the elements in order to be sorted again [26] or reversing the output for the next pass [27]. In particular, the $2$-reverse-pass sortable permutation class has a reasonably small basis, namely $\{2413,2431,23154\}=\{1302,1320,12043\}$ that may prove tractable.

### v6

@@ 5 Open Problems
And the cases restricting pop stacks to particular depths in an analogous way to  Elder (2006); Elder and Goh (2018, 2021); Elder et al. (2015) and/or the expansion in allowable pushes analogous to  Atkinson (1998) can be interesting for words and inversion sequences.
@@
Another avenue of exploration is to consider other sorting machines or algorithms. For instance, the authors considered a sorting algorithm on a single stack prioritizing outputting the next correct entry of the sequence until no other moves were available and then either returning the elements in order to be sorted again Mansour et al. (2019a) or reversing the output for the next pass Mansour et al. (2019b). In particular, the $2$-reverse-pass sortable permutation class has a reasonably small basis, namely $\{2413,2431,23154\}=\{1302,1320,12043\}$ that may prove tractable.
@@
Acknowledgements.
The authors are grateful for the helpful comments from the referees.